Jason's body metabolizes caffeine at a rate of 14.9% per hour (so the amount of caffeine in Jason's body decreases by 14.9% each hour).
a. If Jason consumes a cup of coffee with 84 mg of caffeine in it, how long will it take for Jason's body to metabolize half of the 84 mg of caffeine?
b. If Jason consumes an energy drink with 180 mg of caffeine in it, how long will it take for Jason's body to metabolize half of the 180 mg of caffeine?
c. If Jason consumes a cup of coffee with c mg of caffeine in it, how long will it take for Jason's body to metabolize half of the c mg of caffeine? (Hint: your answer will be a numerical value.)

Answer by Boreal(15235)

84(0.851)^t=41
0.851^t=1/2
ln both sides
t ln 0.851=-0.693
t=4.3 hours
Same answer for b. The half-life time is the same, and the fraction is the same, even if the number of mg changes.
It will take 4.3 hours